The Escobar Corporation uses activity-based costing. They have three cost pools, as follows.

Activity Cost Pool Polishing Finishing Other
Activity Driver Machine Hours Batch Starts None
Budgeted Cost Pool $100,000 $55,000 $112,000
Budgeted Activity Driver 1,600 500 n/a

What are Escobar Corporation's activity rates for these activity cost pools?

a.

Polishing: $62.50 per machine hour; Finishing: $110 per batch start; Other: no rate

b.

Polishing: $100,000 per machine hour; Finishing: $55,000 per batch start; Other: $112,000

c.

Polishing: $166.88 per machine hour; Finishing: $534 per batch start; Other: no rate

d.

Polishing: $0.016 per machine hour; Finishing: $0.01 per batch start; Other: no rate
Answer #1

Activity rate=Respective cost/Respective activity

Activity rate for:

Polishing=(100,000/1600)=$62.5 per machine hour

Finishing=(55000/500)=$110 per batch start

Other=No rate.

Hence the correct option is:

a.Polishing: $62.50 per machine hour; Finishing: $110 per batch start; Other: no rate